# SDP bounds on quantum codes

### v1

@@ 4. Complete SDP hierarchy bounding quantum codes
Our aim is now to express the conditions for the existence of quantum codes as a state polynomial optimization problem. For this we formulate the Knill-Laflamme conditions, the quantum MacWilliams identity, and the constraints for the state to be proportional to a projector of rank $K$ as state polynomial constraints. This gives rise to a complete hierarchy on the existence of an $(\!(n,K,\delta)\!)_{2}$ code.
@@ 4.1. Framework
Consider the $n$-qubit error basis ${\mathcal{E}_{n}}$ formed by tensor products of Pauli matrices [see Eq. (7)]. Define a set of letters $\mathscr{E}_{n}$ so that to each Pauli error $E\in{\mathcal{E}_{n}}$ we associate a letter $e\in\mathscr{E}_{n}$. We require this to be a group homomorphism,
@@
In particular, this implies that $e_{a}^{*}e_{a}=1$ where the empty word $1$ corresponds to identity operator $\mathds{1}$, and for the Pauli group it holds that $\omega_{abc}\in\{\pm 1,\pm i\}$.
Let $M^{(\ell)}$ be the moment matrix from the state polynomial optimization framework [see Eq. (29)], indexed by all state monomials of degree at most $\ell$. We impose all constraints from Eq. (39) on $M^{(\ell)}$. Recall that the entries of the state moment matrix are denoted by $M^{(\ell)}_{vw}={\langle v^{*}w\rangle}$. In order to simplify the notation, we define
(40) $M^{(\ell)}({\langle v^{*}w\rangle}):=M^{(\ell)}_{vw}\,.$
@@ 4.2. Quantum MacWilliams identity
The quantum weight enumerator $A_{j}(\varrho)$ in Eq. (10) is a state polynomial in $\varrho$. We can then define an approximation to the quantum weight enumerator as
@@
so to correspond to $A_{j}(\varrho)=\sum_{E\in{\mathcal{E}_{n}}\,,\operatorname{wt}(E)=j}\operatorname{tr}(E^{\dagger}\varrho)\operatorname{tr}(E\varrho)$. However, the quantum weight enumerator $B_{j}(\varrho)=\sum_{E\in{\mathcal{E}_{n}}\,,\operatorname{wt}(E)=j}\operatorname{tr}(\varrho E^{{\dagger}}\varrho E)$ does not have such a straightforward interpretation as a state polynomial. But recall that the quantum MacWilliams identity [see Eq. (14)] linearly transforms the $A_{i}(\varrho)$ to the $B_{j}(\varrho)$ enumerators. We can then linearly transform $A^{(\ell)}_{j}(\varrho)$ to $B^{(\ell)}_{j}(\varrho)$ using
@@
where $K_{j}(i;n)$ are the Krawtchouk polynomials defined in Eq. (16). Therefore, $B^{(\ell)}_{j}$ provides an approximation to the quantum weight enumerator $B_{j}(\varrho)$.
Note that the $A^{(\ell)}_{j}$ are not necessary compatible with a quantum state, i.e., in general there does not need to be a $\varrho$ such that $A^{(\ell)}_{j}=A_{j}(\varrho)$. However, the state polynomial optimization hierarchy provides a sequence $\{A_{j}^{(\ell)}\}_{\ell=1}^{\infty}$ whose limit converges to an enumerator $A_{j}(\varrho)$ and as a consequence, the sequence $\{B_{j}^{(\ell)}\}_{\ell=1}^{\infty}$ also converges to $B_{j}(\varrho)$.
@@ 4.3. Knill-Laflamme conditions
We now impose the Knill-Laflamme conditions as constraints on $M^{(\ell)}$. These conditions can be formulated in terms of quantum weight enumerators. Likewise, we demand that
@@
Here $S^{(\ell)}_{j}$ is an approximation of the shadow enumerator obtained by Eq. (18). These extra constraints do not have an impact on completeness.
@@ 4.4. Projector constraints
We now impose conditions for a state $\varrho$ to be proportional to a projector of rank $K$ on $M^{(\ell)}$. The state $\varrho$ then fulfills for all $m\in\mathds{N}$,
@@
The converse also holds, namely that a state $\varrho$ satisfying Eq. (45) for all $m\in\mathds{N}$ must be proportional to a projector of rank $K$. To see this, note that the set of $\{\operatorname{tr}(A^{m})\,:\,m=1,\dots,r\}$ determines the characteristic polynomial of a square matrix $A$ of size $r$. The zeros of this characteristic polynomial then determine the eigenvalues of $A$. Fixing the moments as in Eq. (45) implies that the state $\varrho$ is proportional to a projector of rank $K$.
@@
(49) $\sum_{E_{1},\dots,E_{m-1}\in{P}_{n}}\operatorname{tr}(E^{\dagger}_{1}\varrho)\operatorname{tr}(E_{1}E^{\dagger}_{2}\varrho)\dots\operatorname{tr}(E_{m-2}E^{\dagger}_{m-1}\varrho)\operatorname{tr}(E_{m-1}\varrho)=\frac{2^{n(m-1)}}{K^{m-1}}\,.$
Denote by $\mathbf{e}=(e_{1},\dots,e_{m})\in(\mathscr{E}_{n})^{m}$ vectors of length $m$ with entries in $\mathscr{E}_{n}$. Define the state monomial
(50) ${Q}_{(1,\dots,m)}(\mathbf{e})=\begin{cases}{\langle e_{1}\rangle}{\langle e^{*}_{1}e_{2}\rangle}\dots\langle e^{*}_{{m-2}}e_{m-1}\rangle{\langle e^{*}_{m-1}\rangle}\quad&\text{if}\quad m\geq 2\,,\\ {\langle 1\rangle}&\text{if}\quad m=1\,.\end{cases}$
Then, the projector constraints from Eq. (45) can be expressed in terms of state polynomials as the condition that for all $m\in\mathds{N}$,
(51) $\sum_{\textbf{e}\in(\mathscr{E}_{n})^{m}}{Q}_{(1,\dots,m)}(\mathbf{e})=\frac{2^{n(m-1)}}{K^{m-1}}\,.$
One can generalize the monomial of Eq. (50) to all elements of $S_{2\ell}$.
Decompose $\sigma\in S_{2\ell}$ into cycle permutations including one-cycles such that,
(52) $\sigma=(\alpha_{1},\dots,\alpha_{r})\dots(\xi_{1},\dots,\xi_{t})\,.$
Note that $r+\dots+t=2\ell$ since $\sigma\in S_{2\ell}$. Denote by $|\sigma|$ the number of cycles of $\sigma$, including cycles of length $1$. Defining
(53) $P^{(\ell)}_{\sigma}:=\sum_{\mathbf{e_{r}}\in(\mathscr{E}_{n})^{r},\dots,\mathbf{e_{t}}\in(\mathscr{E}_{n})^{t}}\!\!\!\!\!M^{(\ell)}(Q_{(\alpha_{1},\dots,\alpha_{r})}(\mathbf{e_{r}})\dots Q_{(\xi_{1},\dots,\xi_{t})}(\mathbf{e_{t}}))\,,$
we can then impose the constraints from Eq. (53) on $M^{(\ell)}$ as
(54) $P^{(\ell)}_{\sigma}=\Big{(}\frac{2^{n}}{K}\Big{)}^{(r-1)+\dots+(t-1)}=\Big{(}\frac{2^{n}}{K}\Big{)}^{(2\ell-|\sigma|)}\,.$
Consider a $n$-qubit quantum state $\varrho$ and the set of Pauli operators $\mathcal{E}_{n}$, and construct a moment matrix $M^{(\ell)}_{vw}=\operatorname{tr}(E_{v}^{\dagger}E_{w}\varrho)$ indexed by all state monomials in $\varrho$ of at most degree $\ell$. Then $P^{(\ell)}_{\sigma}=\operatorname{tr}(\varrho^{r})\cdots\operatorname{tr}(\varrho^{t})$ for all $\sigma\in S_{2\ell}$, $\ell\geq 2^{n}$ are necessary and sufficient conditions that $\varrho=\Pi/K$ with $\Pi$ a projector of rank $K$ [see Eq. (45)].
@@ 4.5. An SDP hierarchy for qubit codes
We now formulate a semidefinite programming hierarchy that contains the necessary and sufficient conditions for a state to correspond to a quantum code: the first condition is the projector constraints that guarantee that the state is proportional to a projector. The second condition is the Knill-Laflamme conditions, which ensure that this projector corresponds to a subspace with error correction capabilities.
Consider the following hierarchy indexed by $\ell\in\mathds{N}$:
@@
subject to $M^{(\ell)}_{vw}=M^{(\ell)}_{\alpha\beta}\quad\text{if}\quad{\langle v^{*}w\rangle}={\langle\alpha^{*}\beta\rangle}\,,$
$M^{(\ell)}(1)=1\,,$
$KB^{(\ell)}_{j}=A^{(\ell)}_{j}\quad\text{for}\quad 0\leq j<\delta,$
$P^{(\ell)}_{\sigma}=\Big{(}\frac{2^{n}}{K}\Big{)}^{2\ell-|\sigma|}\quad\text{for all}\,\,\,\sigma\in S_{2\ell},$
(55) $A^{(\ell)}$, $B^{(\ell)}_{j}$, $P^{(\ell)}_{\sigma}$ are given by Eqs. (41), (42), (53) respectively.
For every $\ell\in\mathds{N}$, the SDP (4.5) is a relaxation of the optimization problem in Eq. (2.1). This hierarchy allows us to state the following theorem.
@@
($\Longrightarrow$): Let $\varrho$ be a state proportional to the projector of a quantum code with parameters $(\!(n,K,\delta)\!)_{2}$. Given the set of Pauli operators $\mathcal{E}_{n}$, we can construct a moment matrix $M^{(\ell)}_{vw}=\operatorname{tr}(E_{v}^{\dagger}E_{w}\varrho)$ indexed by all state monomials in $\varrho$ of at most degree $\ell$.
It is clear that for all $\ell\in\mathds{N}$, $M^{(\ell)}\succeq 0$ and $M^{(\ell)}(\mathds{1})=1$ [see Section 3.6]. Additionally, $M$ satisfies the projector conditions in Eq. (54) as $\varrho$ is a density matrix proportional to a projector of rank $K$. Finally, $M$ satisfies the state polynomial quantum MacWilliams identity (42) and Knill-Laflamme conditions in enumerator form (43). Therefore, if there exists a code $(\!(n,K,\delta)\!)_{2}$, every level of the hierarchy is feasible.
@@
This state $\varrho$ has three desired properties: First, $\varrho$ is an $n$-qubit state. Recall that, in general, state polynomial optimization optimizes over all compatible Hilbert spaces, operators, and states. In our setting however, Corollary 4 restricts the set of operators to the direct sum of tensor products of Pauli matrices. To see this, consider the GNS construction from the sequence $\{M^{(\ell)}\}^{\infty}_{\ell=1}$: given the sequence, it constructs a set of operators and a state satisfying all relations of the optimization problem. In particular, the constructed operators form a quasi-Clifford algebra over $\mathds{C}$, whose elements square to the identity. Corollary 4 then implies that these operators are isomorphic to the $n$-qubit Pauli basis $\mathcal{E}_{n}$. As a consequence, $\varrho$ is a $n$-qubit state.
Second, $\varrho$ has to be proportional to a projector $\Pi$ of rank $K$. The projector conditions from Eq. (54) are included in the hierarchy via $P^{(\ell)}_{\sigma}$. Since the operators are in the limit isomorphic to the $n$-qubit Pauli basis, the sequence $\{P^{(\ell)}_{\sigma}\}^{\infty}_{\ell=1}$ converges to $\operatorname{tr}(\varrho^{r})\cdots\operatorname{tr}(\varrho^{t})$ with $\varrho$ being a $n$-qubit state. Eq. (54) are now necessary and sufficient conditions for $\varrho$ be proportional to the projector $\Pi$ of rank $K$.
Third, the projector $\Pi=\varrho K$ necessarily satisfies the Knill-Laflamme conditions [see Theorem 1]. Their approximation is included in the hierarchy as $KB_{j}^{(\ell)}=A_{j}^{(\ell)}$ for all $0\leq j<\delta$, becoming exact in the limit $\ell\to\infty$. Similarly to $P^{(\ell)}_{\sigma}$, the sequence $\{A_{j}^{(\ell)}\}_{\ell=1}^{\infty}$ from Eq. (41) converges to the enumerator $A_{j}(\tfrac{\Pi}{K})$. The quantum MacWilliams identity [see Eq. (42)] linearly relates $A^{(\ell)}_{j}$ with $B^{(\ell)}_{j}$, so that the sequence $\{B_{j}^{(\ell)}\}_{\ell=1}^{\infty}$ also converges to $B_{j}(\tfrac{\Pi}{K})$. As a consequence, the projector $\Pi=K\varrho$ fulfills the Knill-Laflamme conditions in the limit $\ell\to\infty$.
Therefore, if every level of the hierarchy is feasible, there exists a state proportional to a projector $\Pi$ on $(\mathds{C}^{2})^{\otimes n}$ corresponding to a quantum code subspace with parameters $(\!(n,K,\delta)\!)_{2}$. This ends the proof. ∎
In the following, we show how an intermediate level of this hierarchy yields a Lovász bound [Section 5], whose symmetrization results a quantum Delsarte bound [Section 7]. Using the Terwilliger algebra of the Hamming association scheme, we also provide efficiently computable symmetry-reductions [Section 8.3]. Finally, we provide an alternative nonexistence proof of the $(\!(7,1,4)\!)_{2}$ code [Section 9.1].

### v2

@@ 4. Complete SDP hierarchy bounding quantum codes
Our aim is now to express the conditions for the existence of quantum codes as a state polynomial optimization problem. For this we formulate the Knill-Laflamme conditions, the quantum MacWilliams identity, and the constraints for the state to be proportional to a projector of rank $K$ as state polynomial constraints. This gives rise to a complete hierarchy for the existence of an $(\!(n,K,\delta)\!)_{2}$ code.
@@ 4.1. Framework
Consider the $n$-qubit error basis $\mathcal{E}_{n}$} formed by $n$-fold tensor products of Pauli matrices [see Eq. (7)]. Define a set of letters $\mathscr{E}_{n}$ so that to each Pauli error $E_{a}\in{\mathcal{E}_{n}}$ we associate a letter $e_{a}\in\mathscr{E}_{n}$. We require this to be a group isomorphism, so that
@@
For the Pauli group it holds that $\omega_{abc}\in\{\pm 1,\pm i\}$. The Pauli relations also imply that $e_{a}^{*}=e_{a}$ and $e_{a}^{*}e_{a}=1$, where the empty word $1$ corresponds to identity operator $\mathds{1}$. For state polynomials, we extend the relations in Eq. (40) linearly to expectations, so that additionally we also have $\langle e_{a}e_{b}\rangle=\omega_{abc}\langle e_{c}\rangle$ if $E_{a}E_{b}=\omega_{abc}E_{c}$. In effect, this imposes the Pauli relations onto the hierarchy in terms of the quotient ring $\mathbb{S}_{n}/I$ where $\mathbb{S}_{n}$ is the set of all state polynomials formed by the letters in $\mathscr{E}_{n}$ and II is the ideal generated by the $n$-qubit Pauli relations. Let $M^{(\ell)}$ be the moment matrix from the state polynomial optimization framework [see Eq. (34)], indexed by all state monomials of degree at most $\ell$. We impose all constraints from Eq. (40) on $M^{(\ell)}$. Recall that the entries of the state moment matrix are denoted by
(41) $M^{(\ell)}({\langle s^{*}t\rangle}):=M^{(\ell)}_{st}\,$.
@@ 4.2. Quantum MacWilliams identity
The quantum weight enumerator $A_{j}(\varrho)$ in Eq. (10) is a state polynomial in $\varrho$. We can then define a state polynomial representing the quantum weight enumerator,
@@
so to correspond to $A_{j}(\varrho)=\sum_{E\in{\mathcal{E}_{n}}\,,\operatorname{wt}(E)=j}\operatorname{tr}(E^{\dagger}\varrho)\operatorname{tr}(E\varrho)$.
However, the quantum weight enumerator $B_{j}(\varrho)=\sum_{E\in{\mathcal{E}_{n}}\,,\operatorname{wt}(E)=j}\operatorname{tr}(\varrho E^{{\dagger}}\varrho E)$ does not have such a straightforward interpretation as a state polynomial. But recall that the quantum MacWilliams identity [see Eq. (14)] linearly transforms the $A_{i}(\varrho)$ to the $B_{j}(\varrho)$ enumerators. We can then linearly transform $A^{(\ell)}_{j}(\varrho)$ to $B^{(\ell)}_{j}(\varrho)$ using
@@
where $K_{j}(i;n)$ are the Krawtchouk polynomials defined in Eq. (16).
@@ 4.3. Knill-Laflamme conditions
We now impose the Knill-Laflamme conditions as constraints on $M^{(\ell)}$. These conditions can be formulated in terms of quantum weight enumerators. Thus we demand that
@@
Here $S^{(\ell)}_{j}$ mirrors the shadow enumerator of Eq. (18). The condition $A^{(\ell)}_{j},\,B^{(\ell)}_{j}\geq 0$ is redundant, since $\Gamma\succeq 0$ contains the elements of $A^{(\ell)}_{j}$ in its diagonal. These extra constraints do not have an impact on completeness.
@@ 4.4. Projector constraints
We now impose conditions for a state $\varrho$ to be proportional to a projector $\Pi$ of rank $K$ on $M^{(\ell)}$. The state $\varrho$ then fulfills for all $m\in\mathds{N}$,
@@
The converse also holds, namely that a state $\varrho$ satisfying Eq. (46) for all $m\in\mathds{N}$ must be proportional to a projector of rank $K$. To see this, note that the set $\{\operatorname{tr}(A^{m})\,:\,m=1,\dots,r\}$ determines the characteristic polynomial of a square matrix $A$ of size $r$. The zeros of this characteristic polynomial then determine the eigenvalues of $A$. Fixing the moments as in Eq. (46) implies that the state $\varrho$ is proportional to a projector of rank $K$.
@@
(50) $\sum_{E_{1},\dots,E_{m-1}\in{\mathcal{E}}_{n}}\operatorname{tr}(E^{\dagger}_{1}\varrho)\operatorname{tr}(E_{1}E^{\dagger}_{2}\varrho)\dots\operatorname{tr}(E_{m-2}E^{\dagger}_{m-1}\varrho)\operatorname{tr}(E_{m-1}\varrho)=\Big{(}\frac{2^{n}}{K}\Big{)}^{m-1}\,.$
The condition (46) for a state to be proportional to a projector of rank $K$, i.e. $\varrho=\Pi/K$, corresponds to the following constraint on the moment matrix for $m=2,\dots,2\ell$.
(51) $P^{(\ell)}_{m}=\sum_{e_{1},\dots,e_{{m-1}}\in{\mathscr{E}_{n}}}M^{(\ell)}\big{(}\langle e^{*}_{1}\rangle\langle e_{1}e^{*}_{2}\rangle\dots\langle e_{{m-2}}e^{*}_{{m-1}}\rangle\langle e_{{m-1}}\rangle\big{)}=\Big{(}\frac{2^{n}}{K}\Big{)}^{m-1}\,.$
Remark 6.
As pointed out to us by Dion Gijswijt, alternatively to Eq. (46), the constraints
(52) $\operatorname{tr}(E_{a}\varrho^{2})=\frac{1}{K}\operatorname{tr}(E_{a}\varrho)\quad\text{for}\quad a=0,\dots,4^{n}-1$
also give necessary and sufficient conditions for $\varrho$ to be proportional to a projector of rank $K$, due to the fact that the $E_{a}$ form a complete basis. These conditions can also be written in terms of state polynomials: similar to Eq. (4.4), expand
$\operatorname{tr}(E_{a}\varrho^{2})$ $=\operatorname{tr}\left((1,2)E_{a}\varrho\otimes\varrho\right)$
(53) $=\frac{1}{2^{n}}\sum_{E_{b}\in\mathcal{E}_{n}}\operatorname{tr}(E_{a}E_{b}^{\dagger}\varrho)\operatorname{tr}(E_{b}\varrho)\,.$
Then, the condition from Eq. (52) can be written in terms of elements of $M^{(\ell)}$ as
(54) $\frac{1}{2^{n}}\sum_{e_{b}\in\mathscr{E}_{n}}M^{(\ell)}\left({\langle e_{a}e_{b}^{\dagger}\rangle}{\langle e_{b}\rangle}\right)=\frac{1}{K}M^{(\ell)}\left({\langle e_{a}\rangle}\right)\quad\text{for}\quad a=0,\dots,4^{n}-1$
The main advantage of Eq. (54) in comparison to Eq. (52) is that already at the level $\ell=2$ a full set of projector conditions is included.
@@ 4.5. An SDP hierarchy for qubit codes
We now formulate a semidefinite programming hierarchy that contains the necessary and sufficient conditions for a state to correspond to a quantum code: the projector constraints guarantee that the state is proportional to a projector of rank $K$, while the Knill-Laflamme conditions ensure that this projector corresponds to a quantum code subspace of minimum distance $\delta$.
From now on consider state polynomials in the quotient ring $\mathbb{S}_{n}/I$, where $I$ is the ideal generated by the $n$-qubit Pauli relations. Then consider the following hierarchy indexed by $\ell\in\mathds{N}$, with the $M^{(\ell)}$ is indexed by all state monomials in $\mathbb{S}_{n}/I$ of degree at most $\ell$ in hermitian variables in $\mathcal{E}_{n}$.
@@
subject to $M^{(\ell)}(1)=1\,$,
$M^{(\ell)}_{st}=M^{(\ell)}_{\alpha\beta}\quad\quad\quad\,\,\,\,\text{if}\quad{\langle s^{*}t\rangle}={\langle\alpha^{*}\beta\rangle}\,$,
$KB^{(\ell)}_{j}=A^{(\ell)}_{j}\quad\quad\quad\,\,\text{for}\quad 0\leq j<\delta,$
$P^{(\ell)}_{m}=\Big{(}\frac{2^{n}}{K}\Big{)}^{m-1}\quad\quad\text{for }\,\,\,m=2,\dots,2\ell\,$,
(55) $A_{j}^{(\ell)}$, $B^{(\ell)}_{j}$, $P^{(\ell)}_{m}$ are given by Eqs. (42), (43), (51) respectively.
For every $\ell\in\mathds{N}$, the SDP (4.5) is a relaxation of the optimization problem in Eq. (2.1).
@@
($\Longrightarrow$): Let $\varrho$ be a state proportional to the projector of a quantum code with parameters $(\!(n,K,\delta)\!)_{2}$. Given the set of Pauli operators $\mathcal{E}_{n}$, we can construct a moment matrix $M^{(\ell)}_{ST}=\operatorname{tr}(S^{\dagger}T\varrho)$ indexed by all Pauli state monomials of at most degree $\ell$. It is clear that for all $\ell\in\mathds{N}$, $M^{(\ell)}\succeq 0$ and $M^{(\ell)}(\mathds{1})=1$ [see Section 3.7]. Additionally, $M$ satisfies the projector conditions in Eq. (51) as $\varrho$ is a density matrix proportional to a projector of rank $K$. Finally, $M$ satisfies the state polynomial quantum MacWilliams identity [Eq. (43)] and Knill-Laflamme conditions in enumerator form [Eq.  (44)]. Thus, if there exists a code with parameters $(\!(n,K,\delta)\!)_{2}$, then every level of the hierarchy is feasible.
@@
find $\mathcal{H},\{X_{a}\}_{a=0}^{4^{n}-1},\varrho$
subject to $KB_{j}(\varrho)=A_{j}(\varrho)\quad\text{for}\quad 0\leq j<\delta\,$,
$A_{j}(\varrho)=\sum_{{X_{a},\,\operatorname{wt}(X_{a})=j}}\operatorname{tr}(X_{a}\varrho)\operatorname{tr}(X_{a}^{\dagger}\varrho)\,$,
$B_{j}(\varrho)=2^{-n}\sum^{n}_{i=0}K_{j}(i;n)A_{i}(\varrho)\,$,
$\text{$\{X_{a}\}_{a=0}^{4^{n}-1}$ satisfies all $n$-qubit Pauli relations}\,$,
$P_{m}=\sum_{{a_{1}},\dots,{a_{m-1}}=0}^{4^{n}-1}\operatorname{tr}(X^{\dagger}_{a_{1}}\varrho)\operatorname{tr}(X_{a_{1}}X^{\dagger}_{a_{2}}\varrho)\dots\operatorname{tr}(X_{a_{m-2}}X^{\dagger}_{a_{m-1}}\varrho)\operatorname{tr}(X_{a_{m-1}}\varrho)$
(56) $=\Big{(}\frac{2^{n}}{K}\Big{)}^{m-1}\quad\text{for all }m\in\mathds{N}\,$,
where $\operatorname{wt}$ is a suitable function defined on the $X_{i}$ corresponding to the weights of the elements in $\mathcal{E}_{n}$, and the $X_{i}$ are hermitian.. This is because the set of constraints from (4.5) is balanced and $4^{n}-\sum_{e\in\mathcal{E}_{n}}ee^{*}=0$.
Now we show that the optimization problem in Eq. (4.5) completely characterizes a $n$-qubit quantum code. First, every finite group algebra decomposes by the Artin-Wedderburn theorem into a direct sum of matrix algebras. In our case, Theorem 3 restricts the set of operators $X_{a}$ to be isomorphic to a direct sum of tensor products of Pauli matrices: considering the commutation relations alone, the operators form a quasi-Clifford algebra over $\mathds{C}$, whose elements square to the identity, and no other relations. The representation in Eq. (38) together with the $n$-qubit Pauli relations baked into $\mathscr{E}_{n}/I$ implies that these operators are isomorphic to the $n$-qubit Pauli basis $\mathcal{E}_{n}$. As a consequence, $\varrho$ is isomorphic to a $n$-qubit quantum state. Finally, the Knill-Laflamme and projector condiditions force $\varrho$ to be the projector onto a code subspace. This ends the proof. ∎
In the following, we show how an intermediate level of this hierarchy yields a Lovász bound [Section 5], whose symmetrization results a quantum Delsarte bound [Section 7]. Using the Terwilliger algebra of the Hamming association scheme, we also provide efficiently computable symmetry-reduced SDP relaxations [Section 8.3]. Finally, we provide an alternative nonexistence proof of the $(\!(7,1,4)\!)_{2}$ code [Section 9.1], and show that $(\!(8,9,3)\!)_{2}$ and $(\!(10,5,4)\!)_{2}$ codes do not exist [Section 9.2].